TWO SHORT PRESENTATIONS FOR LYONS’ SPORADIC SIMPLE GROUP

VOLKER GEBHARDT

Abstract. Using an algorithm developed by the author, two new presentations for R. Lyons’ sporadic simple group are established, which contain fewer relations and are shorter than previously known ones.